one cleaning fluid contains 6% ammonia, while a second contains 26% ammonia. In blending these fluids, at most how many liters of the second may be added to 10 liters of the first, if the percent of ammonia in the mixture is not to exceed 16%?
Found 2 solutions by josgarithmetic, TimothyLamb:
Answer by josgarithmetic(39617)

You can use 16% as the target value. The description is for two materials to blended, and the amount of one is known, but the amount of the other is unknown.

Look at http://www.algebra.com/my/Two-Part-Mixture-with-one-material-quantity-unknown.lesson?content_action=show_dev

Answer by TimothyLamb(4379)

---
6x + 26y = 16(x + y)
x = 10
---
put the system of linear equations into standard form:
---
6x + 26y = 16x + 16y
x = 10
---
10x - 10y = 0
x = 10
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 10 liters of 6% solution
y = at most 10 liters of 26% solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php